If a is acceleration, v is velocity, x is position, and t is time, then
check the validity (wrong or correct) of the following equations using
dimensional analysis:
a) t2 = 2x/a
b) t = x/v
c) a = v/x
d) v = a/t

ALSO ,
The term 1/2 PV^2
rv2 occurs in Bernoulli’s equation in Chapter 15,
with P being the density of a fluid and v its speed. Find the dimensions of
1/2 PV^2

Thank you in advance .
Answer #1
on the first, I will be happy to critique your thinking.

on the second

Pressure*velocity^2=force/area * distance^2/time^2=

force/time^2
answered by: <3Medic+Droid<3
Answer #2
Q2)
(M/L^3) x (L^2/T^2) = M/LT = ML^-1T^-2
answered by: Angel007
Answer #3
in the first i chose last one
but in second p mean density so the formula should be p=m/v m mass , v volume is not it ?